Use the method of substitution to solve the system of linear equations.
2x-y=4
9x-4y=21
The solution is ______.
Please show work!
Thank you
This what I have tried thus far:
2x-y=4
y=2x-4
9x-4y=21
9x-4(2x-4)=21

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

So keep going.

Answer by stanbon(75887)

2x-y=4
y=2x-4
9x-4y=21
9x-4(2x-4)=21
------
9x-8x+16 = 21
x = 5
----
Xolve for "y":
y = 2x-4
y = 2*5-4 = 6
=======================
Cheers,
Stan H.